OpenStudy (anonymous):

Given the function f(x) = 6x^2 −13, what is f(-3)?

OpenStudy (anonymous):

Plug in -3 for your x and solve

OpenStudy (anonymous):

-67?

OpenStudy (anonymous):

f(-3)=6(-3)^2-13 f(-3)=6(9)-13 f(-3)=54-13 f(-3)=41